OpenStudy (anonymous):

If I am evaluating the integral and I find x=3sec(theta) and there is an x^3 on top of the integral. Does this become 3sec^3(theta) or 9sec^3(theta)?

OpenStudy (anonymous):

anyone?

OpenStudy (amistre64):

whats the original integral?

OpenStudy (anonymous):

One second

OpenStudy (amistre64):

x=(3sec(t)) then x^3 = (3sec(t))^3

OpenStudy (anonymous):

\[\int\limits_{}^{}x^3 / \sqrt{x^2+9}\]

OpenStudy (amistre64):

x = 3tan(t) then

OpenStudy (amistre64):

tan^2 + 1 = sec^2

OpenStudy (anonymous):

so do let x=3tan(X) so I get dx = 3sec^2XdX right?

OpenStudy (amistre64):

x = 3tan dx = 3sec^2 dt x^2 = 9tan^3

OpenStudy (anonymous):

So the top is (3tanX)^3? and the bottom is 9(sex^2) then mutlipled by 3sex^2XdX?

OpenStudy (amistre64):

id try to keep the sex out of it :)

OpenStudy (anonymous):

ha good point.

OpenStudy (anonymous):

So it reduces a lot? Will it be 27/9 so be ((3 tan^3(x))/(sec^2X))3sec^2XdX

OpenStudy (amistre64):

\[\int \frac{9tan^3(t)}{3sec(t)}3sec^2(t)dt\] \[\int 9tan^3(t)sec(t)dt\]

OpenStudy (anonymous):

so the answer is 9tan^3X?

OpenStudy (anonymous):

I think the other sec(t) goes away so its just \[\int\limits_{}^{}9\tan^3(t)dt\]

OpenStudy (amistre64):

might have to rename tan^3 into (tan^2)(tan); or simply (sec^2-1)(tan)

OpenStudy (anonymous):

Yeah I think I need to do the first? not really sure never done it ha. Professor likes to give us things we havent done before

OpenStudy (amistre64):

the other sec doesnt just "go away" :) \[9\int (sec^2(t)-1)(sec(t)tan(t))dt\] \[9\int sec^2(t)(sec(t)(tan(t)))-sec(t)tan(t)\ dt\] \[9sec^3(t)-9\int sec(t)tan(t)\ dt\] \[\frac{9}{3}sec^3(t)-9sec(t)+C\]

OpenStudy (amistre64):

then of course you rename your trig back into its appropriate x parts

OpenStudy (anonymous):

Im not sure but I thought the other sec was reduced?

OpenStudy (anonymous):

because dx=3sec^2(t) it reduces the 9(sec^2(t))

OpenStudy (amistre64):

we came up with this after the trig sub \[\int \frac{9tan^3(t)}{\cancel{3sec(t)}}\cancel{3}sec^{\cancel{2}}(t)dt\] \[\int 9tan^3(t)sec(t)dt\]

OpenStudy (anonymous):

So I got my trig function wrong? I thought 1+Tan^2(t)=sec^2(t)?

OpenStudy (amistre64):

\[\sqrt{(3tan(t))^2+9}\to \sqrt{9tan^2(t)+9}\to\sqrt{9}\sqrt{tan^2(t)+1}\to 3sec^2(t)\]

OpenStudy (amistre64):

ugh, typos are pretty lol

OpenStudy (amistre64):

\[...\sqrt{9}\sqrt{tan^2(t)+1}\to 3sec(t)\]

OpenStudy (anonymous):

you are confusing me.. So I was right?

OpenStudy (anonymous):

Ha. So it equals 3sec(t) not 3sec^2(t)?

OpenStudy (amistre64):

right; which cancels out the 3 sec(t) from 3sec^2(t); leaving us up top with just: sec(t)

OpenStudy (anonymous):

Wait... oh tan^2(t)+1 does equal 3sec^2(t) but we have to take the square root?

OpenStudy (anonymous):

I get it now I was just missing the damn square root

OpenStudy (amistre64):

yep, trigs a pain lol

OpenStudy (anonymous):

thanks